What is the velocity at the midway point of a ball able to reach a height y when thrown with upward with an initial velocity v0?


1
Expert's answer
2021-04-05T11:10:58-0400

Let's first find the time that the ball takes to reach the maximum height:


"v=v_0-gt,""0=v_0-gt,""t=\dfrac{v_0}{g}."

Then, we can find the maximum height reached by the ball:


"H=v_0t-\dfrac{1}{2}gt^2,""H=v_0\dfrac{v_0}{g}-\dfrac{1}{2}g(\dfrac{v_0}{g})^2=\dfrac{v_0^2}{2g}."

Let's find the time that the ball takes to reach the midpoint:


"v=v_0-gt_1,""t_1=\dfrac{v_0-v}{g}."

Then, we can find the midpoint height reached by the ball:


"h=v_0t_1-\dfrac{1}{2}gt_1^2,""h=v_0\dfrac{v_0-v}{g}-\dfrac{1}{2}g(\dfrac{v_0-v}{g})^2,""h=\dfrac{v_0^2-v^2}{2g}."

Since "h=\dfrac{H}{2}", we can write:


"\dfrac{v_0^2-v^2}{2g}=\dfrac{v_0^2}{4g},""2v_0^2-2v^2=v_0^2,""v=\dfrac{v_0}{\sqrt{2}}."